# Post-adiabatic waveform-generation framework for asymmetric precessing binaries

### v2

@@ Appendix C Osculating action angles
The functions $q^{r}(\psi^{r},\pi_{i})$ and $q^{\theta}(\psi^{\theta},\pi_{i})$ are given in Eqs. (20b) and (21b) of Ref. [152]. The functions $t_{r}(q_{r},\pi_{i})$ and $t_{\theta}(q_{r},\pi_{i})$ are given in Eqs. (28) and (39) of Ref. [85], with the quantities ‘$\lambda^{(r)}$’ and ‘$\lambda^{(\theta)}$’ therein replaced by $q^{r}/\Upsilon^{r}(\pi_{i})$ and $q^{\theta}/\Upsilon^{\theta}(\pi_{i})$, respectively. Here $\Upsilon^{\alpha}$ are the ‘Mino time’ orbital frequencies [85, 76].
@@
$\frac{d\pi_{i}}{dt}$ $=\varepsilon g_{i}^{(0)}(\mathring{\psi}^{j}_{(0)},\varpi_{j},\tilde{\psi}_{s})$ (262)
@@
$\mathring{\psi}^{i}_{(0)}(\psi^{j},\pi_{j})=\mathring{\psi}^{i}_{(0)}(\psi^{j}_{(0)},\mathring{\pi}_{j})+\varepsilon\left(\delta\psi^{j}\frac{\partial\mathring{\psi}^{i}_{(0)}}{\partial\psi^{j}_{(0)}}+\delta\pi_{j}\frac{\partial\mathring{\psi}^{i}_{(0)}}{\partial\mathring{\pi}_{j}}\right)\\ +\mathcal{O}(\varepsilon^{2}).$ (268)
@@ D.1 Evolution of energy and angular momentum
^$\hat{\nabla}_{\lambda}\hat{v}_{\gamma}=\hat{\nabla}_{[\lambda}\hat{v}_{\gamma]}+\hat{\nabla}_{(\lambda}\hat{v}_{\gamma)}=\hat{\nabla}_{[\lambda}\hat{v}_{\gamma]}+\frac{1}{2}\mathcal{L}_{\hat{v}}\hat{g}_{\lambda\gamma},$ (276)
@@
where we have eliminated terms proportional to ^$\hat{\nabla}_{[\lambda}\hat{\nabla}_{\gamma}v_{\beta]}$ by virtue of the first Bianchi identity. Putting these two terms together we have
@@ D.2 Evolution of the spin-corrected Carter constant
$\frac{d\hat{K}}{d\tau}=V_{1}+V_{2}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3},s^{2}),$ (298)
@@
and used $\mathcal{O}(\varepsilon^{2}s^{0})$ to denote spin-independent $\mathcal{O}(\varepsilon^{2})$ terms. At this stage we can see the evolution equation has no explicit dependence on $\delta S^{\alpha}$ because it would contribute $\mathcal{O}(\varepsilon^{3})$ to $V_{2}$. Like in the analysis of the previous section, the effect of the force generated by $\delta S^{\alpha}$ has been put in a form that does not require an explicit calculation of the nutation angles.
@@
$=\left\langle V_{1}+V_{2}\right\rangle_{\tau}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3},s^{2}),$ (300)
@@
$=-u^{\alpha}u^{\gamma}\left(2h^{\mathrm{R}}_{\delta\alpha;\gamma}-h^{\mathrm{R}}_{\alpha\gamma;\delta}\right)n^{\delta}+\mathcal{O}(\varepsilon^{2}s^{0}),$ (301)
@@
$\delta\Gamma^{\alpha}_{\beta\gamma}\equiv\frac{1}{2}g^{\alpha\delta}\left(2h^{\mathrm{R}}_{\delta(\beta;\gamma)}-h^{\mathrm{R}}_{\beta\gamma;\delta}\right)+\mathcal{O}(\varepsilon^{2}s^{0}).$ (302)
@@
$\delta{}^{*}V_{\alpha\beta\gamma\delta}=\frac{1}{2}\epsilon_{\alpha\beta}{}^{\rho\sigma}\delta V_{\rho\sigma\gamma\delta}+\frac{1}{2}\delta\epsilon_{\alpha\beta}{}^{\rho\sigma}V_{\rho\sigma\gamma\delta}+\mathcal{O}(\varepsilon^{2}s^{0})$ (303)
@@
$-\delta\Gamma^{\rho}_{\delta\gamma}L_{\alpha\beta\rho}-2K_{\gamma\rho}\delta\Gamma^{\rho}{}_{\delta[\beta;\alpha]}+\mathcal{O}(\varepsilon^{2}s^{0}),$ (304)
@@
$\delta\epsilon_{\alpha\beta}{}^{\rho\sigma}=\frac{1}{2}\epsilon_{\alpha\beta}{}^{\rho\sigma}g^{\mu\nu}h^{\mathrm{R}}_{\mu\nu}+2\epsilon_{\alpha\beta}{}^{\mu[\rho}h^{\mathrm{R}\,\sigma]}{}_{\mu}+\mathcal{O}(\varepsilon^{2}s^{0}).$ (305)
@@
$V_{2}=-m_{2}\Bigl{(}4\epsilon_{\alpha\beta}{}^{\mu[\rho}h^{\mathrm{R}\,\sigma]}{}_{\mu}V_{\rho\sigma\gamma\delta}\\ +\epsilon_{\alpha\beta}{}^{\rho\sigma}\delta V_{\rho\sigma\gamma\delta}\Bigr{)}S^{\alpha}u^{\beta}u^{\gamma}u^{\delta}+\mathcal{O}(\varepsilon^{2}s^{0}),$ (306)

### v3

@@ Appendix C Osculating action angles
The functions $q^{r}(\psi^{r},\pi_{i})$ and $q^{\theta}(\psi^{\theta},\pi_{i})$ are given in Eqs. (20b) and (21b) of Ref. [154]. The functions $t_{r}(q_{r},\pi_{i})$ and $t_{\theta}(q_{r},\pi_{i})$ are given in Eqs. (28) and (39) of Ref. [87], with the quantities $\lambda^{(r)}$’ and $\lambda^{(\theta)}$’ therein replaced by $q^{r}/\Upsilon^{r}(\pi_{i})$ and $q^{\theta}/\Upsilon^{\theta}(\pi_{i})$, respectively. Here $\Upsilon^{\alpha}$ are the ‘Mino time’ orbital frequencies [87, 78].
@@
$\frac{d\pi_{i}}{dt}$ $=\varepsilon g_{i}^{(0)}(\mathring{\psi}^{j}_{(0)},\varpi_{j},\tilde{\psi}_{s})$
@@
$\hskip-10.0pt\mathring{\psi}^{i}_{(0)}(\psi^{j},\pi_{j})=\mathring{\psi}^{i}_{(0)}(\psi^{j}_{(0)},\mathring{\pi}_{j})+\varepsilon\left(\delta\psi^{j}\frac{\partial\mathring{\psi}^{i}_{(0)}}{\partial\psi^{j}_{(0)}}+\delta\pi_{j}\frac{\partial\mathring{\psi}^{i}_{(0)}}{\partial\mathring{\pi}_{j}}\right)\\ +\mathcal{O}(\varepsilon^{2}).$ (267)
@@ D.1 Evolution of energy and angular momentum
$\hat{\nabla}_{\lambda}\hat{v}_{\gamma}=\hat{\nabla}_{[\lambda}\hat{v}_{\gamma]}+\hat{\nabla}_{(\lambda}\hat{v}_{\gamma)}=\hat{\nabla}_{[\lambda}\hat{v}_{\gamma]}+\frac{1}{2}\mathcal{L}_{\hat{v}}\hat{g}_{\lambda\gamma},$ (275)
@@
where we have eliminated terms proportional to $\hat{\nabla}_{[\lambda}\hat{\nabla}_{\gamma}v_{\beta]}$ by virtue of the first Bianchi identity. Putting these two terms together we have
@@ D.2 Evolution of the spin-corrected Carter constant
$\frac{d\hat{K}}{d\tau}=V_{1}+V_{2}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3},s^{2}),$ (297)
@@
and used $\mathcal{O}(\varepsilon^{2}\bcancel{s})$ to denote spin-independent $\mathcal{O}(\varepsilon^{2})$ terms. At this stage we can see the evolution equation has no explicit dependence on $\delta S^{\alpha}$ because it would contribute $\mathcal{O}(\varepsilon^{3})$ to $V_{2}$. Like in the analysis of the previous section, the effect of the force generated by $\delta S^{\alpha}$ has been put in a form that does not require an explicit calculation of the nutation angles.
@@
$=\left\langle V_{1}+V_{2}\right\rangle_{\tau}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3},s^{2}),$ (299)
@@
$=-u^{\alpha}u^{\gamma}\left(2h^{\mathrm{R}}_{\delta\alpha;\gamma}-h^{\mathrm{R}}_{\alpha\gamma;\delta}\right)n^{\delta}+\mathcal{O}(\varepsilon^{2}\bcancel{s}),$ (300)
@@
$\delta\Gamma^{\alpha}_{\beta\gamma}\equiv\frac{1}{2}g^{\alpha\delta}\left(2h^{\mathrm{R}}_{\delta(\beta;\gamma)}-h^{\mathrm{R}}_{\beta\gamma;\delta}\right)+\mathcal{O}(\varepsilon^{2}\bcancel{s}).$ (301)
@@
$\delta{}^{*}V_{\alpha\beta\gamma\delta}=\frac{1}{2}\epsilon_{\alpha\beta}{}^{\rho\sigma}\delta V_{\rho\sigma\gamma\delta}+\frac{1}{2}\delta\epsilon_{\alpha\beta}{}^{\rho\sigma}V_{\rho\sigma\gamma\delta}+\mathcal{O}(\varepsilon^{2}\bcancel{s})$ (302)
@@
$-\delta\Gamma^{\rho}_{\delta\gamma}L_{\alpha\beta\rho}-2K_{\gamma\rho}\delta\Gamma^{\rho}{}_{\delta[\beta;\alpha]}+\mathcal{O}(\varepsilon^{2}\bcancel{s}),$ (303)
@@
$\delta\epsilon_{\alpha\beta}{}^{\rho\sigma}=\frac{1}{2}\epsilon_{\alpha\beta}{}^{\rho\sigma}g^{\mu\nu}h^{\mathrm{R}}_{\mu\nu}+2\epsilon_{\alpha\beta}{}^{\mu[\rho}h^{\mathrm{R}\,\sigma]}{}_{\mu}+\mathcal{O}(\varepsilon^{2}\bcancel{s}).$ (304)
@@
$V_{2}=-m_{2}\Bigl(4\epsilon_{\alpha\beta}{}^{\mu[\rho}h^{\mathrm{R}\,\sigma]}{}_{\mu}V_{\rho\sigma\gamma\delta}\\ +\epsilon_{\alpha\beta}{}^{\rho\sigma}\delta V_{\rho\sigma\gamma\delta}\Bigr)S^{\alpha}u^{\beta}u^{\gamma}u^{\delta}+\mathcal{O}(\varepsilon^{2}\bcancel{s}),$ (305)